# Regular Kerr black holes: The fellowship of the string

### v1

@@ Abstract
In their interesting article entitled "Kerrr" black hole: The lord of the string [Phys. Lett. B 688, 82 (2010)], Smailagic and Spallucci report on a regular Kerr-like black hole spacetime with a string in the central region. The metric is of the Kerr-Schild form with a particular mass function, that is chosen in order to avoid the ring singularity of the Kerr metric. However, as it is known, even for a class of well-behaved mass functions, the curvature scalars present a discontinuity in the equatorial plane at the ring. The authors associate such a discontinuity with the presence of a string of matter and join the interior and exterior regions of the string by making a particular choice of the orientations of the normal vectors at each side of the ring. Differently from what was interpreted by the authors, we show here that the string is not rotating in the scenario of that work. We also show that the choice regarding the normal vector orientation leads to a geometry that is not compatible with the center of a rotating black hole in an asymptotically flat spacetime. We then consider all the four possible combinations of the normal vector orientation at each side of the ring and, by using the Darmois-Israel formalism, interpret each resulting solution. We show that, out of the four possibilities, only one of the four models for the thin shell solution at the ring yields the appropriate geometry. More importantly, this solution does not correspond to the one obtained by Smailagic and Spallucci. Finally, we also discuss a scenario where the string is indeed rotating.
@@ I Introduction
The investigation of regular rotating black holes is motivated by the complications within the interior of the Kerr geometry [1], which, while providing a realistic description of the exterior of a rotating black hole, presents issues such as a ring singularity, causality violations, and closed timelike curves within the event horizon [2]. A potential remedy involves replacing the problematic interior region with a regular matter source, mirroring the approach taken for static regular black holes.
@@
In Ref. [3], the authors demonstrated that, in the context of Kerr-Schild spacetimes [4], a Lorentzian coordinate system can be selected in such a way that it leads to the linearization of the Einstein field equations. As an application of this result, the authors showed that complex translations can be applied in order to obtain new solutions for spinning systems, including new anisotropic interior metrics that may be matched to the Kerr metric on an oblate spheroid. This procedure paved the way for constructing rotating counterparts of static and spherically symmetric regular black hole solutions resembling the Kerr metric in the asymptotic region, but with the constant mass parameter replaced by a mass function $m(r)$, so modifying the Kerr geometry close to the central ring. Subsequent explorations have delved into understanding the source generating the rotating metric [5, 6] and a substantial body of work has emerged on regular rotating objects [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 22, 21, 23, 24, 25].
@@
In contrast to the Kerr metric, it was shown by Torres [15] and Maeda [21] that, if the mass function $m(r)$ can be expanded around $r=0$ as $m(r)\approx m_{0}r^{3+\alpha}$ with $m_{0}\neq 0$ and $\alpha\geq 0$, the ring $\mathcal{S}\!\!:\!(r=0,\,\theta=\pi/2)$ corresponds to a conical singularity when extended to the region $r<0$, and not to a scalar polynomial curvature singularity. Moreover, if the extension to the region $r<0$ is not performed, Torres [26] showed that the ring is devoid of conical singularities as well.
@@
Even though the different types of singularities can be avoided in the ring for the Gürses-Gürsey rotating metric, it is well known that the curvature scalars are not well defined there, once they present a finite jump which depends on the path taken when approaching the ring for $\alpha=0$ [7, 8, 15, 21, 19, 24, 25]. On the other hand, if $\alpha>0$ the curvature scalars are continuous [26]. This feature exhibited by the Gürses-Gürsey metric, for $\alpha=0$, has been interpreted by Smailagic and Spallucci [7] as a string substituting the ring singularity observed in the Kerr solution. More specifically, since the limit $r\to 0$ for $\theta=\pi/2$ corresponds to the curvature of a de Sitter-like geometry, while the limit $r\to 0$ for $\theta\neq\pi/2$ corresponds to the curvature of a Minkowski-like geometry [15], the authors glued a Minkowski-like spacetime with a de Sitter-like spacetime at the ring and interpreted the discontinuity in the extrinsic curvature as due to the presence of a string of matter for a unconventional choice of the signs of the normal vectors at ring.
@@
In this work we explore all possible choices of the orientations of the normal vectors at ring. Let us remember that the normal vector of a thin shell plays an important role, since it the determines extrinsic curvature of the shell regarding the spacetime in which the shell is immersed [27]. In fact, to uncover all possible thin shell solutions, it is necessary to take into account that the normal to a shell can have two distinct orientations relative to the center of the coordinates of both the interior and exterior spacetimes (or spacetime regions), even though the normal vector is fixed to point always from a given spacetime to the other [28, 29]. Moreover, we also show that, in the scenario presented in Ref. [7], the string replacing the Kerr ring is actually static. To address this issue, we construct a scenario in which the string is indeed rotating.
II The regular rotating black hole
@@ II.1 The spacetime metric
where $m(r)$ stands for the mass function, $a$ indicates the rotation parameter, and $\Sigma(r,\theta)$ and $\Delta(r)$ are defined by
@@
with $m_{0}\neq 0$ and $\alpha\geq 0$. This assumption guarantees that the desired properties of metric (1) as summarized in Sec. I remain valid, it assures particularly that the curvature scalars do not diverge in the limit $r\to 0$. This is the main reason why such an asymptotic behavior of the mass function is present in most of the works cited here, specially in the work by Smailagic and Spallucci [7].
@@
Hence, for $\theta\neq\pi/2$, the limit $r\to 0$ gives $\mathcal{R}=0$ for all $\alpha\geq 0$. In turn, for $\theta=\pi/2$, Eq. (6) furnishes $2\left(3+\alpha\right)\left(4+\alpha\right)m_{0}r^{\alpha}$ and, therefore, the limit $r\to 0$ gives $\mathcal{R}=0$ for $\alpha>0$, but it gives $\mathcal{R}=24m_{0}$ for $\alpha=0$. In conclusion, if the mass function is of the form $m(r)\sim r^{3}$ close to $r\to 0$, then the Ricci scalar is not well defined there. The same also happens with the other independent curvatures scalars [15].
@@
As just mentioned above, the singular locus of the Kerr metric parameterized by ($r=0,\,\theta=\pi/2$) in the Boyer-Lindquist coordinates can be regularized by replacing the constant mass parameter of the Kerr metric with a power-law mass function of the form (4), but just for $\alpha>0$. For $\alpha<0$ , the locus is still a curvature singularity as in the Kerr metric. The $\alpha=0$ case is special, since the curvature scalars are all finite but present a discontinuity at that locus. This feature motivated the authors of Ref. [7] to introduce a string replacing the ring singularity of the Kerr spacetime. Following this idea, in the present work we explore the proposal by Smailagic and Spallucci in more detail.
@@ II.2 Inside the ring: The disk
Let us notice that the limit $r\to 0$ for $\theta\neq\pi/2$ in Eq. (1) corresponds to a disk $\mathcal{D}$ of radius $a$ that may be parameterized by the polar coordinates $(\rho,\,\varphi)$ such that $0\leq\rho\leq a$ and $0\leq\varphi\leq 2\pi$. The coordinate $\rho$ is given in terms of the Boyer-Lindquist coordinates by
@@ II.3 The ring
The ring $\mathcal{S}$ is at the boundary of the disk defined in the last paragraph, i.e., it may be parameterized by $(\rho=a,0\leq\varphi\leq 2\pi)$. At the ring $\mathcal{S}\!:(r=0,\theta=\pi/2)$, metric (1) reduces to
@@
where, from Eq. (7), $r=0$ and $\theta=\pi/2$ implies $\rho=a$.
@@ II.4 Outside the ring: The equatorial plane
Approaching the ring, but remaining outside, i.e, in the region $\mathcal{A}\!:\!(r\to 0,\,\theta=\pi/2)$, we have $m(\varrho)\approx m_{0}(\varrho^{2}-a^{2})^{3/2}$ and, to the fist order in $\rho^{2}-a^{2}$, the metric given by Eq. (11) reduces to
@@
with $\varrho\in[a,\,a+\varepsilon)$ for $\varepsilon>0$.
@@
In this approximation as made in [7], the terms containing $m_{0}a$ and $m_{0}a^{2}$ that appear in (12) have been ignored and, as a consequence, the exterior metric results a static (nonrotating) metric. In fact, metric (13) may correspond to a de Sitter (dS) or an anti-de Sitter-like (AdS) region, depending on the sign of $m_{0}$, with the curvature scalars being constant. The authors in [7] have chosen $m_{0}>0$, what leads to a de Sitter-like region with positive curvature scalars.

### v2

# Regular Kerr black holes: Junction conditions and the matter content across the ring

@@ Abstract
Regular rotating black holes are usually described by a metric of the Kerr-Schild form with a particular mass function that is chosen to avoid the ring singularity of the Kerr metric and which approaches the Kerr metric at the asymptotic limit. However, as is well known, even for a class of well-behaved mass functions, the curvature scalars present a discontinuity in the equatorial plane at the ring. This discontinuity has been associated with the presence of a string of matter that joins the interior and exterior regions along the equatorial plane. By using the Darmois-Israel junction conditions, we analyze all four possible combinations of the normal vector orientations on each side of the ring, construct the complete stress-energy momentum tensor of the string, and interpret each resulting solution. We show that, out of the four possibilities, only one of the four models for the string solution at the ring yields the appropriate asymptotic geometry. In such a case, the string bears a fluid with nonzero pressure, but with a vanishing line energy density, and it does not rotate at all. Finally, taking an appropriate metric for the exterior region, we also discuss a different scenario in which the matter source at the ring is a rotating lightlike fluid.
@@ I Introduction
The Kerr metric [1] has been widely studied in the general relativity literature for two main reasons. The first reason is because it describes the final state of rotating matter distributions after the complete gravitational collapse, the results of which being rotating black holes. The second reason is related to the intricate properties of such a metric in the region inside the event horizon, which contains a ring singularity, causality violations, and closed timelike curves [2]. A potential remedy for these issues involves replacing the problematic interior region of the Kerr geometry with an appropriate matter source. In addition to avoid the singularity, the source must respect reasonable properties such as stationarity, axial symmetry, separability of the Hamilton-Jacobi equations, and constraining the amount of exotic matter that violates the energy conditions in the region inside the event horizon, once such exotic matter is not observed in astrophysical scenarios [3].
@@
In Ref. [4], the authors demonstrated that, in the context of Kerr-Schild spacetimes [5], a Lorentzian coordinate system can be selected in such a way that it leads to the linearization of the Einstein field equations. As an application of this result, the authors showed that complex translations can be applied in order to obtain new solutions for spinning systems, including new anisotropic interior metrics that may be matched to the Kerr metric on an oblate spheroid. This procedure paved the way for constructing rotating counterparts of static and spherically symmetric regular black hole solutions resembling the Kerr metric in the asymptotic region, but with the constant mass parameter replaced by a mass function $m(r)$, so modifying the Kerr geometry close to the central ring. We refer to this kind of rotating metric as the Gürses-Gürsey metric. Subsequent explorations have been devoted to understanding the source generating the rotating metric [6, 7] and a substantial body of work on regular rotating objects has emerged [8, 9, 10, 11, 12, 13, 14, 15, 17, 16, 18, 19, 20, 21, 27, 22, 3, 23, 24, 25, 26, 28, 29].
@@
In particular, Torres [18] and Maeda [22] have shown that, in contrast to the Kerr metric, if the mass function $m(r)$ can be expanded around $r=0$ as $m(r)\approx m_{0}r^{3+\alpha}$ with $m_{0}\neq 0$ and $\alpha\geq 0$, the ring $\mathcal{S}\!:\!(r=0,\,\theta=\pi/2)$ corresponds to a conical singularity when extended to the region $r<0$, and not to a scalar polynomial curvature singularity. Moreover, if the extension to the region $r<0$ is not performed, Torres [30] showed that the ring is devoid of conical singularities as well.
Although these different types of singularity can be avoided in the ring for the Gürses-Gürsey rotating metric [4], it is well known that the curvature scalars are not well-defined there, once they present a finite jump which depends on the path taken when approaching the ring for $\alpha=0$ [8, 9, 18, 22, 27, 28, 29]. On the other hand, if $\alpha>0$ the curvature scalars are continuous and does not depend on the path taken when approaching the ring [30]. Interestingly, the curvature scalars are also well-defined and continuous for the mass function with exponential suppression, that is, $m(r)=m_{0}e^{-l/r}$ [15, 3], where $l$ is a parameter that quantifies the deviation from the Kerr geometry. These kinds of regular Kerr geometries with an exponential mass function represent rotating regular black holes with an asymptotically Minkowski core [3].
@@
The feature exhibited by the Gürses-Gürsey metric for $\alpha=0$ at the ring has been interpreted by Smailagic and Spallucci [8] as a string replacing the ring singularity observed in the Kerr solution. More specifically, since the limit $r\to 0$ for $\theta=\pi/2$ corresponds to the curvature of a de Sitter-like geometry, while the limit $r\to 0$ for $\theta\neq\pi/2$ corresponds to the curvature of the Minkowski geometry [18], the authors glued a section of the Minkowski spacetime with a de Sitter-like spacetime at the ring and, for unconventional orientations of the normal vectors, interpreted the discontinuity of the extrinsic curvature as due to the presence of a string of matter at the ring.
@@
The finite jump presented by the curvature scalars at the Kerr ring deserves further investigation and, therefore, motivated by the previous contributions mentioned above, in this work we study in detail the proposal of replacing the ring by a string and explore all possible choices of the signs of the normal vectors at the ring. Let us remind that the normal vector of a thin shell plays an important role, since it determines extrinsic curvature of the shell regarding the spacetime in which the shell is immersed [31]. In fact, to uncover all possible thin shell solutions, it is necessary to take into account that the normal to a shell can have two distinct orientations relative to the center of the coordinates of both the interior and exterior spacetimes (or spacetime regions), even though the normal vector is fixed to point always from a given spacetime to the other [32, 33, 34]. Moreover, we also show that the string that replaces the Kerr ring in the scenario presented in Ref. [8] is actually static. To address this issue, we construct a scenario in which the string is indeed rotating.
The present work is organized as follows. In Sec. II, we review the regular Kerr geometry and explore its main properties, in particular, the behavior of the relevant curvature scalars and other quantities close to the ring. We also present the metrics for the regions inside, outside, and the ring. The Darmois-Israel junction conditions are briefly described in Sec. III. The matching at the ring using the same spacetime construction as in Ref. [8] is done in Sec. IV, where we uncover all possible thin-shell (string) solutions and show that only one of the four possibilities yields the appropriate geometry and show that neither of them corresponds to a rotating string. In Sec. V, we present a new spacetime construction in which the string is indeed rotating, and we analyze the four possible thin-shell solutions. Our final comments are made in Sec. VI.
II The regular rotating geometry
@@ II.1 The spacetime metric
where $m(r)$ stands for the mass function, $a$ indicates the rotation parameter, which we assume to be positive, and $\Sigma(r,\theta)$ and $\Delta(r)$ are defined by
@@
with $m_{0}\neq 0$ and $\alpha\geq 0$. This assumption guarantees that the desired properties of metric (1) as summarized in Sec. I remain valid, it assures particularly that the curvature scalars do not diverge in the limit $r\to 0$. This is the main reason why such an asymptotic behavior of the mass function is present in most of the works cited here, especially the work by Smailagic and Spallucci [8].
@@
Hence, for $\theta\neq\pi/2$, the limit $r\to 0$ gives $\mathcal{R}=0$ for all $\alpha\geq 0$. In turn, for $\theta=\pi/2$, Eq. (6) furnishes $2\left(3+\alpha\right)\left(4+\alpha\right)m_{0}r^{\alpha}$, and therefore, the limit $r\to 0$ gives $\mathcal{R}=0$ for $\alpha>0$, but it gives $\mathcal{R}=24m_{0}$ for $\alpha=0$. In conclusion, if the mass function is of the form $m(r)\sim r^{3+\alpha}$ close to $r\to 0$, then the Ricci scalar is well-defined for all values of $\alpha>0$. However, for $\alpha=0$, in the limit to the ring ($r\to 0,\theta\to\pi/2$) the Ricci scalar is a double-valued function, i.e., $\mathcal{R}(r=0,\theta=\pi/2+\epsilon)=0$ but $\mathcal{R}(r=0,\theta=\pi/2)=24m_{0}$, and hence it is not well-defined there. The same happens also with the other independent curvature scalars [18].
For instance, the Ketschmann scalar results as
$\begin{split}\mathcal{K}=&\dfrac{4\,m_{0}^{2}r^{2(\alpha+2)}}{\Sigma^{2}(r,\theta)}\bigg{[}\dfrac{384\,r^{8}}{\Sigma^{4}(r,\theta)}-\dfrac{192\left(\alpha+6\right)\,r^{6}}{\Sigma^{3}(r,\theta)}\\ &+\dfrac{8\,r^{4}}{\Sigma^{2}(r,\theta)}\big{[}27+2(\alpha+3)(3\alpha+22)\big{]}\\ &-\dfrac{4\,r^{2}}{\Sigma(r,\theta)}\big{[}(\alpha+4)\big{(}2\alpha(\alpha+9)+39\big{)}\big{]}\\ &+{4(\alpha+3)^{2}(\alpha+4)^{2}}\bigg{]}.\end{split}$ (7)
The last equation shows that the behavior of the Kretschmann scalar in the region $r\to 0$ is similar to the Ricci scalar. Assuming $\theta\neq\pi/2$, the limit $r\to 0$ gives ${\cal K}=0$ for $\alpha\geq 0$. However, for $\theta=\pi/2$ we get ${\cal K}(r,\pi/2)_{r\to 0}=4\left(\alpha^{4}+6\alpha^{3}+17\alpha^{2}+28\alpha+24\right)m_{0}^{2}\,r^{2\alpha}$, which gives zero for $\alpha>0$, but gives ${\cal K}=96m_{0}^{2}$ for $\alpha=0$. In conclusion, we verify that for $\alpha=0$ the Kretschmann scalar is double-valued in the Kerr ring, similar to the Ricci scalar, an undesired feature that deserves further investigation.
The source that generates the Gürses-Gürsey geometry is an anistropic fluid [4, 6, 7], whose energy-momentum tensor may be cast into the form
$T^{\mu\nu}=\rho_{m}\,e_{0}^{\ \mu}e_{0}^{\ \nu}+\mathfrak{p}_{1}e_{1}^{\ \mu}e_{1}^{\ \nu}+\mathfrak{p}_{2}e_{2}^{\ \mu}e_{2}^{\ \nu}+\mathfrak{p}_{3}e_{3}^{\ \mu}e_{3}^{\ \nu},$ (8)
[…]
$e_{0}^{\ \mu}=\frac{1}{\sqrt{\pm\Delta(r)\Sigma}}\Big{(}\left[r^{2}+a^{2}\right]\delta^{\mu}_{\ t}+a\,\delta^{\mu}_{\ \varphi}\Big{)},$
$e_{1}^{\ \mu}=\sqrt{\frac{\pm\Delta(r)}{\Sigma}}\delta^{\mu}_{\ r},\ \ e_{2}^{\ \mu}=\frac{1}{\sqrt{\Sigma}}\delta^{\mu}_{\ \theta},$ (9)
$e_{3}^{\ \mu}=\frac{1}{\sqrt{\Sigma}\sin\theta}\big{[}a\sin^{2}\theta\delta^{\mu}_{\ t}+\delta^{\mu}_{\ \varphi}\big{]},$
is the well known Carter’s orthonormal frame, which diagonalizes the energy-momentum tensor according to Eq. (8). For regular black holes, the plus sign applies to the regions outside the event horizon and inside the Cauchy horizon, with $e_{0}^{\ \mu}$ being a unit timelike vector that describes the four-velocity of the fluid. In turn, the minus sign applies to the region between the event horizon and the Cauchy horizon, with the vectors $e_{0}^{\ \mu}$ and $e_{1}^{\ \mu}$ switching roles.
The total energy density $\rho_{m}$, the radial pressure $\mathfrak{p}_{1}$, and the tangential pressures $\mathfrak{p}_{2}$ and $\mathfrak{p}_{3}$ are given by [6]
$\rho_{m}(r,\theta)=-\mathfrak{p}_{1}(r,\theta)=\frac{r^{2}m^{\prime}(r)}{4\pi\Sigma^{2}},$ (10)
$\mathfrak{p}_{2}(r,\theta)=\mathfrak{p}_{3}(r,\theta)=\frac{r^{2}m^{\prime}(r)}{4\pi\Sigma^{2}}-\frac{1}{8\pi\Sigma}\Big{(}r\,m(r)\Big{)}^{\prime\prime},$ (11)
with the prime indicating derivative with respect to the coordinate $r$. The above relations for the fluid quantities hold for any mass function $m(r)$, but we are interested in the cases where the mass function is such that $m(r)\sim r^{3}$ as $r\to 0$.
The timelike vector $e_{0}^{\ \mu}$ is the four-velocity of a stationary fluid that rotates with the angular velocity
$\Omega(r)=\frac{e_{0}^{\ \varphi}}{e_{0}^{\ t}}=\frac{a}{r^{2}+a^{2}}.$ (12)
Another interesting property of the regular Kerr metric is the so-called frame dragging. The frame-dragging velocity $\omega$ is defined by considering an observer that, starting from infinity with zero angular momentum, falls freely toward the region defined by $r=0$. Such conditions define the so-called zero angular momentum observers. In the Boyer-Lindquist coordinates, one has
$\omega(r,\theta)=-\frac{g_{\varphi t}}{g_{\varphi\varphi}}.$ (13)
where
$g_{\varphi t}$ $=-\frac{2r\,m(r)\,a\sin^{2}\theta}{\Sigma},$ (14)
$g_{\varphi\varphi}$ $=\left(r^{2}+a^{2}+\frac{2r\,m(r)\,a^{2}\sin^{2}\theta}{\Sigma}\right)\sin^{2}\theta.$ (15)
The frame-dragging velocity (13) is well-defined everywhere, including the ring ($r=0$, $\theta=\pi/2$) where it vanishes. In the equatorial plane, $\theta=\pi/2$, it gives $\omega(r,\pi/2)=2m(r)\,r^{2}a\big{/}\left[r^{2}\left(r^{2}+a^{2}\right)+2m(r)\,r\,a^{2}\right]$. So, by considering a mass function that close to $r=0$ is of the form (4), the limit $r\to 0$ yields
$\lim_{r\to 0}\omega(r,\pi/2)=0.$ (16)
This important result is different from the singular Kerr metric in which case one has $\lim_{r\to 0}\omega(r,\pi/2)=1/a$. For $\theta\neq\pi/2$, the situation here is the same as for the singular Kerr geometry, namely,
$\lim_{r\to 0}\omega(r,\theta\neq\pi/2)=0,$ (17)
so that the frame-dragging velocity is well-defined everywhere in the regular Kerr-like geometry.
@@
As just mentioned above, the singular locus of the Kerr metric parameterized by ($r=0,\,\theta=\pi/2$) in the Boyer-Lindquist coordinates can be regularized by replacing the constant mass parameter of the Kerr metric with a power-law mass function of the form (4), but just for $\alpha>0$. For $\alpha<0$, the locus is still a curvature singularity as in the Kerr metric. The case $\alpha=0$ is special since the curvature scalars are all finite but present a discontinuity at that locus. This feature motivated the authors of Ref. [8] to introduce a string replacing the ring singularity of the Kerr spacetime. Following this idea, in the next section, we shall explore the proposal by Smailagic and Spallucci in more detail.
@@ II.2 Inside the ring: The disk
The limit $r\to 0$ for $\theta\neq\pi/2$ of the metric (1) corresponds to a disk $\mathcal{D}$ of radius $a$ that can be parameterized by the polar coordinates $(\rho,\,\varphi)$ so that $0\leq\rho\leq a$ and $0\leq\varphi\leq 2\pi$. The coordinate $\rho$ is given in terms of the Boyer-Lindquist coordinates by
@@ II.3 The ring
The ring $\mathcal{S}$ is the boundary of the disk defined in the last paragraph, i.e., it may be parameterized by $(\rho=a,0\leq\varphi\leq 2\pi)$. At the ring $\mathcal{S}\!:\!(r=0,\theta=\pi/2)$, metric (1) reduces to
@@
where, from Eq. (18), the assumptions $r=0$ and $\theta=\pi/2$ imply $\rho=a$.
@@ II.4 Outside the ring: The equatorial plane
Approaching the ring, but remaining outside, i.e., in the region $\mathcal{A}\!:\!(r\to 0^{+},\,\theta=\pi/2)$, the mass function $m(\varrho)$ is given approximately by the expression in Eq. (4), namely,
$m(\varrho)=m_{0}(\varrho^{2}-a^{2})^{3/2},$ (23)
and, to the fist order in $\rho^{2}-a^{2}$, the metric (22) reduces to
@@
with $\varrho\in[a,\,a+\varepsilon)$ for $\varepsilon>0$. This is the metric assumed to describe the exterior region of the ring. Note that (24) is the general rotating Kerr-like metric (1) with $\theta=\pi/2$ and $m(r)=m_{0}r^{3}$, i.e., it is the complete rotating metric (22) with $m(\varrho)=m_{0}\left(\varrho^{2}-a^{2}\right)^{3/2}$, without any further approximation.
@@
In comparison to the rotating metric (24), the metric (25) neglects the terms containing $m_{0}a\left(\varrho^{2}-a^{2}\right)$ and $m_{0}a^{2}\left(\varrho^{2}-a^{2}\right)$. As a consequence, the $g_{t\varphi}$ coefficient vanishes and then the result is a static (nonrotating) metric. In fact, the metric (25) can be considered an approximation for the static de Sitter (dS) metric, or for the anti-de Sitter metric, depending on the sign of $m_{0}$. The authors in [8] have chosen $m_{0}>0$, which leads to a de Sitter-like metric. In turn, metric (24) is appropriate for the exterior metric, since it is exactly the rotating metric (1) in the equatorial with a properly chosen mass function.
It is worth mentioning that most regular black holes described in the literature present a de Sitter-type core, which corresponds to $m_{0}>0$. Even though we are not aware of any specific work that constructs regular black holes with anti-de Sitter-type cores, i.e., with $m_{0}<0$, in Ref. [35] the author shows that, among other kinds of objects, it is possible to construct such regular black holes.